OpenStudy (anonymous):

Find all solutions of the equation in the interval [0,2pi). 2(sin^2)x = 2 + cosx

OpenStudy (mertsj):

replace sin^2x with 1-cos^2x

OpenStudy (anonymous):

Alright, I'm working it out in my notebook hold on

OpenStudy (anonymous):

Do you get 2(cos^2)x -cosx = 0?

OpenStudy (mertsj):

\[2(1-\cos ^2x)=2+\cos x\] \[2-2\cos ^2x=2+\cos x\] \[-2\cos ^2x-\cos x=0\] \[\cos ^2x+\cos x=0\]

OpenStudy (mertsj):

Whoops. Forgot the 2

OpenStudy (mertsj):

\[2\cos ^2x+\cos x=0\]

OpenStudy (mertsj):

Factor out cos x. Set each factor equal to 0 and solve.

OpenStudy (anonymous):

Ah, thank you